OpenStudy (anonymous):

Without dividing, do you expect 7/8 divided by 11/16 to be greater than or less than 1? Explain.

OpenStudy (anonymous):

(7/8)/(11/16), using the chain rule = = (7*16)/(8*11) = 112/88 > 1

OpenStudy (hoblos):

7/8 > 11/16 => 7/8 divided by 11/16 must be greater than 1

OpenStudy (anonymous):

Well, it says without actually dividing. The answer should be as follows. 7/8 is almost 1. 11/16 is significantly below 1. Dividing anything close to one by something significantly below 1 is likely to result in a number greater than 1.

OpenStudy (anonymous):

without dividing, doesnt mean you cant multiply. Use the chain rule as i say. It's the best way for you to make sure